Grade 11Physical Chemistry

Density of dry air containing only N2 and O2 is 1.15g/l at 740mm of hg amd 300k.What is %composition of N2 by mass in the air?

Jai
8 Years agoGrade 11

1 Answer

Abhiraj Pathak
8 Years ago
According to ideal gas law, PM = dRTwhere d is the density, M is the molar mass or in this case average molar mass and P,T and R have obvious parameters.M =dRT/PMTaking R = 0.0821 L atm P =740/760 atmT = 300K,Substituting these values we have M = 29.05Let x be the percentage of N2 in airThus 1-x will be the percentage of O2Thus 28x + (1-x)*32 = 29.054Solving x = 73.6%Kindly approve